OpenStudy (anonymous):

and passing through the given point. Then write the equation in standard form. Slope-3/8 ; through (5, 3) Choose one answer. a. 3x - 8y = 39 b. 3x + 8y = -39 c. 3x + 8y = 39 d. 8x + 3y = -39

hero (hero):

Use the point slope formula: \[y - y_1 = m(x - x_1)\]

hero (hero):

So: m = -3/8 (x_1,y_1 )=(5,3) Plugging in the given values into the formula: y - 3 = -3/8(x - 5)

hero (hero):

y - 3 = -3x/8 + 15/8 y = -3x/8 + 15/8 + 3 y = -3x/8 + 15/8 + 24/8 y = -3x/8 + 39/8 y = (-3x + 39)/8